OpenStudy (anonymous):

{(sin(π/3+Δx)−sin(π/3)} limΔx→0 ________________________ Δx

OpenStudy (lgbasallote):

first, you need to know that sin(A+ B) = sinAcosB + cosAsinB so... sin(pi/3)cos(delta x) + cos(pi/3)sin(delta x) - sin(pi/3) all over delta x sin(pi/3)(cos(delta x) - 1) + cos(pi/3)(sin delta x) / delta x (sqrt 3/2 (cos delta x - 1) + sin delta x/2)/delta x ...im stuck for the moment :)) maybe you'll have an idea here :D

OpenStudy (anonymous):

I'm looking over your work, but i'm not sure where the -1 came from

OpenStudy (accessdenied):

I think, because this is essentially the derivative of sin(x) at x=pi/3, you would just take the derivative of sin(x), which is cos(x), and then use x=pi/3. (The -1 came from the fact that he factored out a sin(pi/3) from the numerator, and the last term was "-sin(pi/3)")